# Discarding Lavrentiev’s Gap in Non-automonous and Non-Convex Variational Problems

### v1

@@ Abstract
under a Dirichlet boundary condition. Here, $\Omega$ is a bounded Lipschitz open set in ${\mathbb{R}^{N}}$, $N\geq 1$ and the function $f$ is required to be measurable with respect to the spacial variable, continuous with respect to the second one, and continuous and comparable to convex with respect to the last variable. Moreover, we assume that $f$ satisfies a natural condition balancing the variations with respect to the first variable and the growth with respect to the last one. Remarkably, typical conditions that are usually imposed on $f$ to discard the Lavrentiev gap are dropped here: we do not require $f$ to be bounded or convex with respect to the second variable, nor impose any condition of $\Delta_{2}$-kind with respect to the last variable.
[…]
@@ 1 Introduction
We deal with the Lavrentiev phenomenon when the infimum of a variational problem over a family of regular functions is strictly greater than the infimum taken over all functions satisfying the same boundary conditions. This study dates back to [Lav]. Understanding the phenomenon in the multidimensional calculus of variations holds significant importance due to its implications not only in mathematical analysis [Buttazzo-Mizel, BonCel, DeATr, bacomi-cv], but also in various applied fields such as physics, engineering, and materials science, where optimization problems frequently arise, cf. [numerics1, numerics2]. By its very nature, excluding the Lavrentiev phenomenon is closely related to the approximation theory [AmGoSe, Bor-Chl, C-b, Dal-Maso, Gossez]. Deepening the study of the phenomenon prompts the development of new methodologies for tackling non-smooth optimization problems. Our main focus is on excluding the Lavrentiev gap between Sobolev and Lipschitz maps for the scalar functional
$E_{\Omega}(u)=\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}\,,$ (1.1)
where $\Omega$ is a bounded Lipschitz open set in ${\mathbb{R}^{N}}$, $N\geq 1$. We assume convexity of $f$ only with respect to the last variable and allow a dependence of $f$ on the spacial variable $x$. In view of counterexamples [Balci23, badi, zh, eslemi, bcdfm], it is natural that we impose on $f$ a condition balancing the variations with respect to the first variable and the growth with respect to the last one. In fact, the absence of the Lavrentiev phenomenon might be inferred if every function from a relevant space might be approximated in a suitable sense, which typically requires this balanced behaviour of $f$ as well, cf. [AmGoSe, C-b, bgs-arma, Bor-Chl, BCM, hahab, Lukas23]. Due to the fact that $f$ is not assumed to be convex with respect to the second variable, the functional $E_{\Omega}$ is not convex either. Generally speaking, non-convex problems introduce additional complexities that lead to discontinuities and non-smoothness of the solutions, making it more difficult to exclude the Lavrentiev phenomenon. Such problems attract broad attention [Bousquet-Pisa, BocMur, Cupini2024]. We follow the ideas of [Bousquet-Pisa] that get rid of the typical structural assumptions, see e.g. [MaTr, DeATr, BonCel, Ekeland-Temam]. The main idea behind our approach relies on the analysis of a convexification of the energy $E_{\Omega}$, which acts on the subgraph of $u$. Ultimately, since the problem that we study is non-autonomous, we shall consider a version of the convexified energy that involves the spacial variable. The major challenge lies in finding a way to control this spacial dependence of the convexified energy knowing how the original energy varies. Besides allowing a fairly general dependence on the first two variables, the key advantages of our method are that, unlike other contributions, we are able to avoid imposing any assumption such as the $\Delta_{2}$-condition on $f(x,t,\cdot)$, nor any structural assumptions on the dependence of $f$ with respect to the last variable. Moreover, we cover the one dimensional case $N=1$ together with higher dimensions within one reasoning.
@@
Our main results state the absence of the Lavrentiev gap between Lipschitz functions and energy space of the functional $E_{\Omega}$, defined in (1.1), under a Dirichlet boundary condition formulated in terms of a Lipschitz function ${\varphi}$. In order to present our framework, let us settle some more notation. We consider the space $W^{1,1}_{\varphi}(\Omega)$ of those functions $u\in W^{1,1}(\Omega)$ which coincide with ${\varphi}$ on $\partial\Omega$ and the subset $W^{1,\infty}_{\varphi}(\Omega)$ of those Lipschitz continuous functions $u$ on $\Omega$ which agree with ${\varphi}$ on $\partial\Omega$. Let us define
@@
For a function $f$ and a given ball $B$, we denote
@@
In order to exclude the Lavrentiev phenomenon for $E_{\Omega}$ from (1.1) we shall assume conditions balancing the growth of $f$ with respect to the last variable with controlled behaviour of $f$ under small spacial perturbations called later ‘anti-jump’ conditions (given by ($H^{\rm iso}_{0}$), ($H_{\Delta_{2}}$), ($H^{\textrm{conv}}$), and ($H^{\textrm{iso}}$)). Let us start with presenting our results in a simplified setting, namely when $f$ is isotropic, i.e., $f(x,t,\xi)=f(x,t,|\xi|)$ or satisfies a typical growth condition of doubling-type. In the case when $f$ is isotropic, our toy model result under no doubling-type condition reads as follows.
@@
Let $\Omega$ be a bounded Lipschitz open set in ${\mathbb{R}^{N}}$, ${\varphi}:{\mathbb{R}^{N}}\to{\mathbb{R}}$ be Lipschitz continuous, $p\geq 1$, and $f:\Omega\times{\mathbb{R}}\times{[0,\infty)}\to{[0,\infty)}$ be a Carathéodory function, which is convex with respect to the last variable and satisfies $f(x,t,0)=0$ for a.e. $x\in\Omega$ and all $t\in{\mathbb{R}}$. Suppose that for every $L=(L_{1},L_{2})\in(0,\infty)^{2}$, there exists a constant $C_{L}>0$ such that for a.e. $x,y\in\Omega$ and every $(t,\xi)\in[-L_{1},L_{1}]\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
@@
We stress that the above result directly generalises [bgs-arma, Theorem 2.3], as for non-trivial $u$-dependence of the integrand (which is possibly unbounded and non-convex), no need for the integrand to satisfy the $\Delta_{2}$-condition nor to have super-linear growth in the last variable (nor to grow faster than a fixed power function). Consequently, we retrieve classical results for variable exponents and double phase functionals [eslemi, zh], as well as new ones for the latter [bcdfm]. We also refer to Theorem 4 below for the extended isotropic and orthotropic version of this result.
The isotropic structure of the functionals considered in Theorem 1 can be relaxed to the fully anisotropic one, i.e., when $f$ depends on the last variable not necessarily via its length. We present here a simple fully anisotropic consequence of our main result. The price for such a clean expression is the need for requiring the typical doubling growth assumption.
Theorem 2.
Suppose $\Omega$ is a bounded Lipschitz open set in ${\mathbb{R}^{N}}$ and ${\varphi}:{\mathbb{R}^{N}}\to{\mathbb{R}}$ is Lipschitz and continuous. Let $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ be a Carathéodory function which is convex with respect to the last variable, which satisfies $f(x,t,0)=0$ and $f(x,t,\cdot)\in\Delta_{2}$ for a.e. $x\in\Omega$ and all $t\in{\mathbb{R}}$. Suppose that for every $L=(L_{1},L_{2})\in(0,\infty)^{2}$, there exists a constant $C_{L}>0$ such that for a.e. $x\in\Omega$ and every $(t,\xi)\in[-L_{1},L_{1}]\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
Consider the functional $E_{\Omega}$ defined in (1.1). Then, for every $u\in{\mathscr{E}_{\varphi}}(\Omega)$ there exists a sequence $(u_{n})_{n\in{\mathbb{N}}}\subset W^{1,\infty}_{\varphi}(\Omega)$ such that $u_{n}\to u$ in $W^{1,1}(\Omega)$ as $n\to\infty$ and
$\lim_{n\to\infty}E_{\Omega}(u_{n})=E_{\Omega}(u)\,.$
Both above results, presented as Theorems 1 and 2, are special instances of Theorem 3, which can be found in Section 2. Let us emphasize the following facts.

### v3

# Discarding Lavrentiev’s Gap in Non-autonomous and Non-Convex Variational Problems

@@ Abstract
under a Dirichlet boundary condition. Here, $\Omega$ is a bounded Lipschitz open set in $\mathbb{R}^{N}$}, $N\geq 1$ and the function $f$ is required to be measurable with respect to the spatial variable, continuous with respect to the second one, and convex with respect to the last variable. Under these assumptions alone, Lavrentiev gaps may occur, as illustrated by classical examples in the literature. We identify an additional natural condition on $f$ to discard such phenomena, that can be interpreted as a balance between the variations with respect to the first variable and the growth with respect to the last one. This unifies most of the structural assumptions that have been introduced so far to prevent the occurence of Lavrentiev gaps. Remarkably, typical assumptions that are usually imposed on $f$ in this setting are dropped here: we do not require $f$ to be bounded or convex with respect to the second variable, nor impose any condition of $\Delta_{2}$-kind with respect to the last variable.
@@ 1 Introduction
We are interested in excluding the Lavrentiev phenomenon, which occurs when the infimum of a variational problem over a family of regular functions is strictly greater than the infimum taken over all functions satisfying the same boundary conditions. Understanding the phenomenon in the multidimensional calculus of variations is of significant importance due to its implications not only in mathematical analysis but also in various applied fields, such as physics, engineering, and materials science, where optimization problems frequently arise, cf. [6, 8, 22, 20, 28, 19, 47] and references therein. By its very nature, excluding the Lavrentiev phenomenon is closely related to the approximation theory [3, 11, 23, 27, 36]. Moreover, it is intrinsically connected to the regularity theory in the Calculus of Variations. Indeed, for a Sobolev minimizer to be Lipschitz continuous, the absence of the Lavrentiev phenomenon is clearly a necessary condition. It turns out that this is also a sufficient condition for the regularity of minimizers for a large family of functionals [4, 32, 25, 24, 40, 39, 10].
Lavrentiev [44] presented the first example of a one-dimensional variational problem for which the infimum on the set of absolutely continuous functions differs from the infimum on the set of Lipschitz functions. This example was later simplified by Manià [46], who considered the following problem:
$I_{M}:u\in W^{1,1}(0,1)\mapsto\int_{0}^{1}(u(x)^{3}-x)^{2}u^{\prime}(x)^{6}\,{\,{\rm d}}x,$
where the admissible functions $u$ are subject to the boundary conditions $u(0)=0$ and $u(1)=1$. Then, one has
$\inf_{\begin{subarray}{c}u\in W^{1,1}(0,1),\\ u(0)=0,u(1)=1\end{subarray}}I_{M}(u)<\inf_{\begin{subarray}{c}u\in W^{1,\infty}(0,1),\\ u(0)=0,u(1)=1\end{subarray}}I_{M}(u)\,.$ (1.1)
This example demonstrates that the Lavrentiev phenomenon is not related to the smoothness of the integrand, as it can occur even when the latter is a polynomial. The example also shows that convexity with respect to the last variable does not discard the Lavrentiev phenomenon. In fact, it persists for modified versions of the above problem where the integrand is uniformly convex in the last variable [9]. One may wonder whether an even more elementary example exists, where the Lagrangian would depend only on $(x,u^{\prime})$ or $(u,u^{\prime})$. In the one-dimensional case, the answer is negative, see [44, 1, 21]. The three variables $(x,u,u^{\prime})$ have to be summoned for the Lavrentiev phenomenon to occur.
In the higher-dimensional case, we study the Lavrentiev gap between Sobolev and Lipschitz maps for a general scalar functional of the form
$E_{\Omega}:u\in W^{1,1}(\Omega)\mapsto\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}\,$, (1.2)
where $\Omega$ is a bounded Lipschitz open set in $\mathbb{R}^{N}$}, $N\geq 1$. Our standing assumptions require that $f$ be a nonnegative Carathéodory map which is convex with respect to the last variable. For integrands $f$ depending just on $(x,\nabla u)$ or just on $(u,\nabla u)$, one may be tempted to conjecture for the case $N\geq 2$ the non-occurrence of the Lavrentiev phenomenon established when $N=1$. It turns out that the situation is much more complex in the higher-dimensional setting. For autonomous Lagrangians, namely those which do not depend on the spatial variable $x$, this conjecture is correct. There is no Lavrentiev phenomenon in that case, as in the one-dimensional situation [16]. On the contrary, when $f$ depends just on $(x,\nabla u)$, Lavrentiev phenomena have been detected for many functionals [49, 32, 7, 13].
Let us conjure the functional which has attracted notable attention in the past few years [12, 10, 25, 24, 13, 18], the double phase functional:
$u\mapsto\int_{\Omega}|\nabla u|^{p}+a(x)|\nabla u|^{q}\,{\,{\rm d}}x\,$, (1.3)
where $1<p<q$ and $a$ is a nonnegative and bounded function. A considerable amount of work has been achieved to obtain sharp conditions on $p,q$ and the behaviour of $a$ that prevent the Lavrentiev phenomenon [32, 7, 13, 12, 18]. Significant effort has been devoted to extending these conditions to more general Lagrangians $f$, which still depend only on $x$ and $\nabla u$. As many examples suggest, the harsh changes of growth of the Lagrangian $f(x,\nabla u)$ with respect to $\nabla u$ when $x$ varies are responsible for the Lavrentiev phenomenon. In the case of the double phase functional (1.3), the Lagrangian has a $p$ growth (in terms of $\nabla u$) at the points $x$ where $a$ vanishes, whereas it has a $q$ growth where $a$ is strictly positive. If the variations of $a$ are not suitably small, in terms of the difference qpq-p or the ratio $q/p$, then the Lavrentiev phenomenon may occur. The competition of several regimes of $\nabla u$-growth associated with several subregions of $\Omega$ is acknowledged as the only source of Lavrentiev phenomena. From a technical point of view, however, most of the articles discarding the Lavrentiev phenomenon have found it convenient to introduce additional growth restrictions in the $\nabla u$ variable, such as the $\Delta_{2}$ condition. For later use, let us define this condition for a general function $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{\mathbb{R}}$: There exists $C>0$ such that for every $(x,t,\xi)\in\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}$, one has
$f(x,t,2\xi)\leq C(f(x,t,\xi)+1).$ (1.4)
We then write $f(x,t,\cdot)\in\Delta_{2}$. Polynomials do satisfy this assumption, which, on the contrary, fail for integrands involving the exponential map. In this paper, our main result does not require such growth conditions.
For more general Lagrangians where the $u$ dependence comes into play, it is a priori unclear whether this competition between different $\nabla u$-growth regimes is solely responsible for the Lavrentiev phenomenon. Let us recall that in the one-dimensional case, the Lavrentiev phenomenon can occur only when $f$ depends on the three variables, emphasizing the role of $u$ in producing the gaps between the two infima, compare (1.1). This suggests that in the higher dimensional case, new phenomena could arise when passing from Lagrangians depending just on $(x,\nabla u)$ to Lagrangians fully depending on the three variables $(x,u,\nabla u)$.
The main goal of this article is to address this possibility. Roughly speaking, our main results state that the spatial variations of the growth with respect to $u$ and with respect to $\nabla u$ should now be taken into account. This does not mean, however, that those variables $u$ and $\nabla u$ have to be considered on an equal footing, for otherwise unnecessary restrictions would be artificially introduced on the $u$-dependence. Without entering into too many details at this stage, let us mention two important aspects:
  • 1.
    The summability of $W^{1,p}$ functions is (much) higher than the summability of their gradients, as a consequence of the Sobolev embeddings. It is thus expectable that the conditions involving the summability exponents on $u$ and $\nabla u$ should not be the same.
  • 2.
    More importantly, $f$ is not assumed to be convex with respect to $(u,\nabla u)$ but just with respect to $\nabla u$.
Regarding the second point, we stress the fact that the convexity assumption just with respect to $\nabla u$ is the natural condition to impose in the general framework of the direct method of existence in the Calculus of Variations. Indeed, this assumption is essentially equivalent to the weak lower semicontinuity of the functional. In contrast, requiring the convexity of $f$ jointly with respect to $u$ and $\nabla u$ is unnecessarily restrictive in this regard.
In fact, the study of the Lavrentiev phenomenon can be placed within a broader framework. The weak sequential lower semicontinuity of the functional $E_{\Omega}$ is equivalent to the identity
$E_{\Omega}(u)=\inf\left\{\liminf_{n\to+\infty}E_{\Omega}(u_{n}):(u_{n})_{n\geq 1}\subset W^{1,1}(\Omega),u_{n}\rightharpoonup_{W^{1,1}}u\right\}.$
The corresponding relaxed functional on $W^{1,\infty}(\Omega)$ is instead defined as
$E^{\textrm{rel}}_{\Omega}(u):=\inf\left\{\liminf_{n\to+\infty}E_{\Omega}(u_{n}):(u_{n})_{n\geq 1}\subset W^{1,\infty}(\Omega),u_{n}\rightharpoonup_{W^{1,1}}u\right\}.$
The problem of the integral representation of $E^{\textrm{rel}}_{\Omega}(u)$ amounts to the existence of some function $g:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{\mathbb{R}}$ such that
$E^{\textrm{rel}}_{\Omega}(u)=\int_{\Omega}g(x,u(x),\nabla u(x))\,dx.$ (1.5)
One may expect that $g=f$ under suitable conditions on $f$. As a consequence of our main results, we can assert that the assumptions on $f$ which discard the Lavrentiev phenomenon are sufficient to establish (1.5) with $g=f$. Actually, we have an even stronger conclusion: there is no Lavrentiev gap, in the sense that for every $u\in W^{1,1}(\Omega)$, there exists a sequence $(u_{n})_{n\geq 1}\subset W^{1,\infty}(\Omega)$ which strongly converges to $u$ in $W^{1,1}(\Omega)$ and such that $\lim_{n\to+\infty}E_{\Omega}(u_{n})=E_{\Omega}(u)$. Moreover, we can ensure that each $u_{n}$ agrees with $u$ on $\partial\Omega$ (provided of course that the trace of $u$ on $\partial\Omega$ is Lipschitz continuous). It is expected, as the absence of the Lavrentiev phenomenon is typically inferred by excluding the Lavrentiev gap, see [3, 23, 18, 11, 14, 38, 43].
@@
Main results. In order to present our framework, let us settle some more notation. We consider the space $W^{1,1}_{\varphi}(\Omega)$ of those functions $u\in W^{1,1}(\Omega)$ which coincide with $\varphi$} on $\partial\Omega$ and the subset $W^{1,\infty}_{\varphi}(\Omega)$ of those Lipschitz continuous functions $u$ on $\Omega$ which agree with $\varphi$} on $\partial\Omega$. Let us define
@@
Our main results state the absence of the Lavrentiev gap between Lipschitz functions and the energy space of the functional $E_{\Omega}$, defined in (1.2), under a Dirichlet boundary condition formulated in terms of a Lipschitz function $\varphi$}. As already mentioned, this requires a condition on $f$ which controls the growth in the last two variables $u$ and $\nabla u$ when the spatial variable $x$ varies. Since this condition, in its more general unifying form (see ($H^{\textrm{conv}}$) below), looks quite uninterpretable at first glance, we prefer to begin with some specific situations where our main result can be formulated in more friendly-looking terms. Let us emphasize, however, that the following statements are already interesting in their own right: they extend and simplify many related theorems on the non-occurrence of the Lavrentiev gap.
In the first simplified situation that we address, we assume that $f$ is isotropic, i.e., $f(x,t,\xi)=f(x,t,|\xi|)$, where $|\xi|$ is the Euclidean norm of $\xi$. In that case, the condition connecting the $(t,\xi)$-growth of $f$ with the spatial variable $x$ requires that:
For every $L=(L_{1},L_{2})\in(0,\infty)^{2}$, there exists a constant $C_{L}>0$ such that for a.e. $x,y\in\Omega$ and every $(t,\xi)\in[-L_{1},L_{1}]\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
Hence, if $x$ and $y$ are not too far apart and $\xi$ is not too large, then $f(x,t,\xi)$ and $f(y,t,\xi)$ are comparable. Observe that the distance between $x$ and $y$ on the one hand, and the Euclidean norm of $\xi$ on the other hand, are controlled with the same parameter $\varepsilon$}. In particular, we do not require that the right-hand side holds for any $\xi$, but just for those which are suitably small with respect to |xy||x-y|. To the best of our knowledge, this type of condition was first introduced by Zhikov in his seminal paper [49] on the double phase problem, compare (1.3). It should be emphasized that the second variable $t$ and the third variable $\xi$ are not on equal terms in ($H^{\rm iso}_{0}$). More specifically, once $L$ is fixed, the variable $t$ is confined to a bounded interval, and the large values of $t$ are thus ignored, whatever the choice of $\varepsilon$} may be. For isotropic Lagrangians, ($H^{\rm iso}_{0}$) is enough to discard any Lavrentiev gap. This is the content of the following statement, which can be deduced from our main result Theorem 3:
Theorem 1.
Let $\Omega$ be a bounded Lipschitz open set in $\mathbb{R}^{N}$}, ${\varphi}:{\mathbb{R}^{N}}\to{\mathbb{R}}$ be Lipschitz continuous, $p\geq 1$, and $f:\Omega\times{\mathbb{R}}\times{[0,\infty)}\to{[0,\infty)}$ be a Carathéodory function, which is convex with respect to the last variable and satisfies $f(x,t,0)=0$ for a.e. $x\in\Omega$ and all $t\in{\mathbb{R}}$. Suppose that ($H^{\rm iso}_{0}$) holds. Consider the functional $E_{\Omega}$ defined in (1.2). Then, for every $u\in{\mathscr{E}_{\varphi}}(\Omega)\cap W^{1,p}(\Omega)$, there exists a sequence $(u_{n})_{n\in{\mathbb{N}}}\subset W^{1,\infty}_{\varphi}(\Omega)$ such that $u_{n}\to u$ in $W^{1,p}(\Omega)$ as $n\to\infty$ and
$\lim_{n\to\infty}E_{\Omega}(u_{n})=E_{\Omega}(u)\,.$
We stress that the above result directly generalises [18, Theorem 2.3], since it includes a non-trivial $u$-dependence of the integrand, which is possibly unbounded and non-convex. Moreover, there is no need here for the integrand to satisfy the $\Delta_{2}$-condition nor to have super-linear growth in the last variable (nor to grow faster than a fixed power function). Consequently, we retrieve classical results for variable exponents and double phase functionals [32, 49], as well as new ones for the latter [13]. We present novel examples of functionals to which Theorem 1 applies in Section 2. We also refer to Theorem 5 below for the extended isotropic and orthotropic version of this result.
The formulation of ($H^{\rm iso}_{0}$) can be simplified by introducing the following notation: For a measurable function $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{\mathbb{R}}$ and a given ball $B\subset{\mathbb{R}^{N}}$ intersecting $\Omega$, we denote
@@
Then, ($H^{\rm iso}_{0}$) reads
$|\xi|\leq L_{2}{\varepsilon}^{-\min\big(1,\frac{N}{p}\big)}\Longrightarrow f(x,t,|\xi|)\leq C_{L}\Big(f_{B(x,{\varepsilon})}^{-}(t,\xi)+1\Big)\,.$
The isotropic structure of the functionals considered in Theorem 1 can be relaxed to the fully anisotropic one, i.e., when $f$ depends on the last variable not necessarily via its length. We present here a simple, fully anisotropic consequence of our main result, Theorem 3. The new condition on $f$ balancing the $(t,\xi)$-growth in terms of the spatial variations now reads:
For every $L=(L_{1},L_{2})\in(0,\infty)^{2}$, there exists a constant $C_{L}>0$ such that for a.e. $x\in\Omega$ and every $(t,\xi)\in[-L_{1},L_{1}]\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
Observe that ($H_{\Delta_{2}}$) differs from ($H^{\rm iso}_{0}$) only through the way that the length of $t$ and $\xi$ are measured. While in ($H^{\rm iso}_{0}$), the quantities $f(x,t,\xi)$ and $f_{B(x,{\varepsilon})}^{-}(t,\xi)$ are comparable when the norm of $\xi$ alone is not too large, in ($H_{\Delta_{2}}$) instead, this holds under a suitable control on $f_{B(x,{\varepsilon})}^{-}(t,\xi)$. As illustrated by many examples in Section 2, the latter condition often yields sharper results, even for isotropic functionals, provided that the $\Delta_{2}$ condition (1.4) is satisfied.
@@
Suppose $\Omega$ is a bounded Lipschitz open set in $\mathbb{R}^{N}$} and ${\varphi}:{\mathbb{R}^{N}}\to{\mathbb{R}}$ is Lipschitz continuous. Let $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ be a Carathéodory function which is convex with respect to the last variable, which satisfies $f(x,t,0)=0$ and $f(x,t,\cdot)\in\Delta_{2}$ for a.e. $x\in\Omega$ and all $t\in{\mathbb{R}}$. Suppose that ($H_{\Delta_{2}}$) holds. Consider the functional $E_{\Omega}$ defined in (1.2). Then, for every $u\in{\mathscr{E}_{\varphi}}(\Omega)$, there exists a sequence $(u_{n})_{n\in{\mathbb{N}}}\subset W^{1,\infty}_{\varphi}(\Omega)$ such that $u_{n}\to u$ in $W^{1,1}(\Omega)$ as $n\to\infty$ and
@@
We shall stress that already this special case is new, as it allows for an essentially non-convex structure of $f$. Indeed, the dependence of $f$ on the second variable is just continuous, but not bounded nor convex as e.g. in [11, 14, 17]. This makes the proof a non-trivial task, which cannot be solved by merely varying previous reasonings. Moreover, let us emphasize the following facts.